# Directional data analysis using the spherical Cauchy and the Poisson kernel-based distribution

### v4

@@ 2 The spherical Cauchy and Poisson kernel-based distributions
$f({\bf y})=C_{d}\left(\frac{1-\rho^{2}}{1+\rho^{2}-2{\bf y}^{\top}{\bf m}}\right)^{d},$ (1)
@@
$f({\bf y})=C_{d}\left(\sqrt{\|\boldsymbol{\mu}\|^{2}+1}-{\bf y}^{\top}\boldsymbol{\mu}\right)^{-d}=C_{d}\left(\sqrt{\gamma^{2}+1}-\alpha\right)^{-d},$ (2)
where $\bm{\mu}\in\mathbb{R}^{d+1}$ is the unconstrained location parameter, $\alpha={\bf y}^{\top}\boldsymbol{\mu}$, $\gamma=\|\boldsymbol{\mu}\|$, ${\bf m}=\boldsymbol{\mu}/\gamma$ and $\rho=(\sqrt{\gamma^{2}+1}-1)/\gamma$ ($\gamma=\frac{2\rho}{1-\rho^{2}}$). The benefit of this parameterization is that the maximisation with respect to the location parameter is unconstrained.
@@
$f({\bf y})=C_{d}\frac{1-\rho^{2}}{\left(1+\rho^{2}-2{\bf y}^{\top}{\bf m}\right)^{(d+1)/2}},$ (3)
@@
$f({\bf y})=C_{d}\left(\sqrt{\gamma^{2}+1}-\alpha\right)^{-(d+1)/2}\left(1-\frac{(\sqrt{\gamma^{2}+1}-1)^{2}}{\gamma^{2}}\right)^{-(d-1)/2}.$ (4)
Figure 1 presents the contour plots for some location parameter and two values of the concentration parameter $\rho$ for the SC distribution. The contour plots have the same shape, but the tails of the PKB decay slower than those of the SC distribution. The density value of the SC will be larger than the density value of the PKB, that is, if ${\bf y}^{\top}\boldsymbol{\mu}<(1+\rho^{2})/2$.
@@ 2.1 Simulation of directional vectors
In order to simulate random directional vectors ${\bf y}_{i}$ from the $\text{SC}({\bf y};{\bf m},\rho)$, Kato and McCullagh, (2020) proposed a simple procedure that relies upon the uniform distribution and then perform some simple calculations. The fact that no rejection sampling is necessary, a strategy common with other distributions, is an appealing one. The two steps of the algorithm are described below.
@@
    Set ${\bf y}_{i}=\frac{\left({\bf u}_{i}+\rho{\bf m}\right)\left(1-\rho^{2}\sum_{j=1}^{d+1}{\bf m}_{j}^{2}\right)}{\sum_{j=1}^{d+1}\left({\bf u}_{i}+\rho{\bf m}_{j}\right)^{2}}+\rho{\bf m}$.
In order to simulate random directional vectors ${\bf y}_{i}$ from the $\text{PKB}({\bf y};{\bf m},\rho)$, Sablica et al., (2023) proposed a rejection sampling that is computationally more expensive than the simulation of the SC distribution.
@@ 2.2 Maximum likelihood estimation
The log-likelihood for a sample of directional vectors ${\bf y}_{i}$, $i=1\ldots,n$ of the SC distribution, using Eq. (2) is given by
@@
where $C_{d}$ denotes the normalizing constant. To perform maximum likelihood estimation (MLE) the Newton-Raphson (NR) algorithm can be employed, to maximise $\ell_{SC}$ (5). The starting value for the NR algorithm is the sample mean vector for which the log-likelihood value is computed. At each successive step the estimated mean vector is updated via $\boldsymbol{\mu}^{t+1}=\boldsymbol{\mu}^{t+1}-{\bf H}^{-1}{\bf J}$ and the algorithm terminates when the difference between two successive log-likelihood values is smaller than some tolerance value $\epsilon$ (for instance $\epsilon=10^{-6}$).
@@
$\ell_{SC}=n\log{C_{d}}+nd\log\left(1-\rho^{2}\right)-d\sum_{i=1}^{n}\log\left(1+\rho^{2}-2\rho{\bf y}_{i}^{\top}{\bf m}\right).$ (6)
@@ 1. Step 1
    Start with an initial mean direction given by $\hat{\bf m}=\frac{\bar{\bf y}}{\|\bar{\bf y}\|}$, where $\bar{\bf y}$ denotes the sample mean vector.
@@ 3. Step 3
    $\ell_{SC}=n\log{C_{d}}+nd\log\left(1-\hat{\rho}^{2}\right)-d\sum_{i=1}^{n}\log\left(1+\hat{\rho}^{2}-2\hat{\rho}{\bf y}_{i}^{\top}{\bf m}\right)+\lambda\left({\bf m}^{\top}{\bf m}-1\right).$ (7)
@@
    $\frac{\partial\ell_{SC}}{\partial{\bf m}}=d\sum_{i=1}^{n}\frac{2\rho{\bf y}_{i}}{1+\hat{\rho}^{2}-2\hat{\rho}{\bf y}_{i}^{\top}{\bf m}}+2\lambda{\bf m}={\bf 0}.$
    The updated mean direction is given by the unit vector parallel to $\sum_{i=1}^{n}\frac{\hat{\rho}{\bf y}_{i}}{1+\hat{\rho}^{2}-2\hat{\rho}{\bf y}_{i}^{\top}\hat{{\bf m}}}$.
@@ 2.3 Log-likelihood ratio test for equality of two location parameters
$\ell_{0}=n_{1}d\log\left(1-\rho_{1}^{2}\right)-d\sum_{i=1}^{n_{1}}\log\left(1+\rho_{1}^{2}-2\rho_{1}{\bf y}_{1i}^{\top}{\bf m}_{c}\right)+n_{2}d\log\left(1-\rho_{2}^{2}\right)-d\sum_{i=1}^{n_{2}}\log\left(1+\rho_{2}^{2}-2\rho_{2}{\bf y}_{2i}^{\top}{\bf m}_{c}\right),$ (9)
where $n_{j}$ and ${\bf y}_{ji}$ refer to the $j$-th sample size and the $i$-th observation of sample $j$, respectively, for $j=1,2$. Under the alternative hypothesis, the two location parameters are not equal and hence the relevant log-likelihood to be maximised is given by
$\ell_{1}=n_{1}d\log\left(1-\rho_{1}^{2}\right)-d\sum_{i=1}^{n_{1}}\log\left(1+\rho_{1}^{2}-2\rho_{1}{\bf y}_{1i}^{\top}{\bf m}_{1}\right)+n_{2}d\log\left(1-\rho_{2}^{2}\right)-d\sum_{i=1}^{n_{2}}\log\left(1+\rho_{2}^{2}-2\rho_{2}{\bf y}_{2i}^{\top}{\bf m}_{2}\right),$ (10)
@@ 2.4 Regression analysis
When a set of $p$ covariates is present, we link the location parameter to the covariates by $\bm{\mu}_{i}={\bf B}^{\top}{\bf x}_{i}$, where ${\bf B}=\left(\bm{\beta}_{1}^{\top},\ldots,\bm{\beta}_{p}^{\top}\right)$ denotes the matrix of the regression coefficients and ${\bf x}_{i}$ denotes a row of the design matrix. The relevant log-likelihood, using (2), becomes
$\ell_{SC}=\log\left(C_{d}\right)-d\sum_{i=1}^{n}\log\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right),$ (11)
where $\alpha_{i}={\bf y}_{i}^{\top}\bm{\mu}_{i}$ and $\gamma_{i}=\|\bm{\mu}_{i}\|$.
The advantage of the parametrization in Eq. (2) is evident in the regression setting. Similarly to Paine et al., (2020) and Tsagris and Alzeley, (2024) the errors are anisotropic, because we do not assume a common concentration parameter. Each directional vector has its own concentration parameter $\gamma_{i}$ that is linked to $\rho_{i}$ as mentioned earlier. The matrix of the regression coefficients is again estimated via the NR algorithm.
@@
where $\alpha_{i}$ and $\gamma_{i}$ are the same as in the case of the SC regression.
@@ 2.5 Discriminant analysis
Furthermore, we assume that we have at hand a labeled directional dataset, that is, for each observation $\textbf{y}_{i}\in\mathbb{S}^{d}$, the allocation variable $c_{i}\in\{1,\ldots,J\}$ is also recorded, $i=1,\ldots,n$. The goal is to assign a feature observation $\textbf{y}_{0}$ into one among the $J$ classes. Under the maximum likelihood discriminant analysis framework (Mardia et al.,, 1979, 301–305), the rule is to allocate a new observation vector ${\bf y}_{0}\in\mathbb{S}^{d}$ in the group whose likelihood is maximized, i.e,
@@
$\log{\left(\sqrt{\|\boldsymbol{\mu}_{j}\|^{2}+1}-{\bf y}^{\top}_{0}\boldsymbol{\mu}_{j}\right)},\ \ j=1,\ldots,J$
and allocate ${\bf y}_{0}$ to the group $j$ that minimizes the above quantity. In practice, these parameters are unknown, so they are estimated, separately, for each group. In the case of the PKB distribution the allocation rule is straightforward to write down analytically, but evidently, in a more complicated form. In this case, ${\bf y}_{0}$ is allocated to the group $j$ with the minimum value in the following quantity
@@ 2.6 Model based clustering
where $f(\cdot;\bm{\mu}_{j})$ denotes the probability density function of the spherical Cauchy or Poisson kernel-based distribution with (unconstrained) parameter $\bm{\mu}_{j}\in\mathbb{R}^{d+1}$ for $j=1,\ldots,K$. Notice that the previous equation in case where the $j$-th element of $\bm{Z}_{i}$ is equal to 1, collapses to $\bm{Y}_{i}|(Z_{ij}=1)\sim f(\cdot;\bm{\mu}_{j})$. It follows that the joint distribution of $\bm{Y}_{i},\bm{Z}_{i}$ is
@@
Initialization of the EM algorithm demands extra care in order to avoid convergence to local maxima. Some indicative works on this topic include Biernacki et al., (2003), Karlis and Xekalaki, (2003), Fraley et al., (2005), Baudry and Celeux, (2015), Papastamoulis et al., (2016), Michael and Melnykov, (2016). In our implementation of the mixtures of SC distributions, the suggested initialization is based on $k$-means starting values since we observed that for this particular class of finite mixture models this strategy gave comparable results to more sophisticated initialization schemes (see Section 2.3 of Papastamoulis, (2023)). For the case of the PKBD we relied on the implementation of Golzy and Markatou, (2020).
@@ 3.2 Hypothesis testing for two location parameters
Table 3: Estimated type I error (nominal error is set to 5%) and estimated power when the data are generated from the SC distribution with concentration parameters equal to $\rho_{1}=0.3$ and $\rho_{2}=0.8$, for the first and second sample, respectively. The angular difference between the two location directions is given by $\theta$.

### v5

@@ 2 The spherical Cauchy and Poisson kernel-based distributions
$f(\bm{y})=C_{d}\left(\frac{1-\rho^{2}}{1+\rho^{2}-2\rho\bm{y}^{\top}{\bf m}}\right)^{d},$ (1)
@@
$f(\bm{y})=C_{d}\left(\sqrt{\|\boldsymbol{\mu}\|^{2}+1}-\bm{y}^{\top}\boldsymbol{\mu}\right)^{-d}=C_{d}\left(\sqrt{\gamma^{2}+1}-\alpha\right)^{-d},$ (2)
where $\bm{\mu}\in\mathbb{R}^{d+1}$ is the unconstrained location parameter, $\alpha=\bm{y}^{\top}\boldsymbol{\mu}$, $\gamma=\|\boldsymbol{\mu}\|$, ${\bf m}=\boldsymbol{\mu}/\gamma$ and $\rho=(\sqrt{\gamma^{2}+1}-1)/\gamma$ ($\gamma=\frac{2\rho}{1-\rho^{2}}$). The benefit of this parameterization is that the maximisation with respect to the location parameter is unconstrained.
@@
$f(\bm{y})=C_{d}\frac{1-\rho^{2}}{\left(1+\rho^{2}-2\rho\bm{y}^{\top}{\bf m}\right)^{(d+1)/2}},$ (3)
@@
$f(\bm{y})=C_{d}\left(\sqrt{\gamma^{2}+1}-\alpha\right)^{-(d+1)/2}\left(1-\frac{(\sqrt{\gamma^{2}+1}-1)^{2}}{\gamma^{2}}\right)^{-(d-1)/2}.$ (4)
Figure 1 presents the contour plots for some location parameter and two values of the concentration parameter $\rho$ for the SC distribution. The contour plots have the same shape, but the tails of the PKB decay slower than those of the SC distribution. The density value of the SC will be larger than the density value of the PKB, that is, if $\bm{y}^{\top}\boldsymbol{\mu}<(1+\rho^{2})/2$.
@@ 2.1 Simulation of directional vectors
In order to simulate random directional vectors $\bm{y}_{i}$ from the $\text{SC}(\bm{y};{\bf m},\rho)$, Kato and McCullagh, (2020) proposed a simple procedure that relies upon the uniform distribution and then perform some simple calculations. The fact that no rejection sampling is necessary, a strategy common with other distributions, is an appealing one. The two steps of the algorithm are described below.
@@
    Set $\bm{y}_{i}=\frac{\left({\bf u}_{i}+\rho{\bf m}\right)\left(1-\rho^{2}\sum_{j=1}^{d+1}{\bf m}_{j}^{2}\right)}{\sum_{j=1}^{d+1}\left({\bf u}_{i}+\rho{\bf m}_{j}\right)^{2}}+\rho{\bf m}$.
In order to simulate random directional vectors $\bm{y}_{i}$ from the $\text{PKB}(\bm{y};{\bf m},\rho)$, Sablica et al., (2023) proposed a rejection sampling that is computationally more expensive than the simulation of the SC distribution.
@@ 2.2 Maximum likelihood estimation
The log-likelihood for a sample of directional vectors $\bm{y}_{i}$, $i=1\ldots,n$ of the SC distribution, using Eq. (2) is given by
@@
where $C_{d}$ denotes the normalizing constant. To perform maximum likelihood estimation (MLE) the Newton-Raphson (NR) algorithm can be employed, to maximise $\ell_{SC}$ (5). The starting value for the NR algorithm is the sample mean vector for which the log-likelihood value is computed. At each successive step the estimated mean vector is updated via $\boldsymbol{\mu}^{t+1}=\boldsymbol{\mu}^{t}-{\bf H}^{-1}{\bf J}$ and the algorithm terminates when the difference between two successive log-likelihood values is smaller than some tolerance value $\epsilon$ (for instance $\epsilon=10^{-6}$).
@@
$\ell_{SC}=n\log{C_{d}}+nd\log\left(1-\rho^{2}\right)-d\sum_{i=1}^{n}\log\left(1+\rho^{2}-2\rho\bm{y}_{i}^{\top}{\bf m}\right).$ (6)
@@ 1. Step 1
    Start with an initial mean direction given by $\hat{\bf m}=\frac{\bar{\bm{y}}}{\|\bar{\bm{y}}\|}$, where $\bar{\bm{y}}$ denotes the sample mean vector.
@@ 3. Step 3
    $\ell_{SC}=n\log{C_{d}}+nd\log\left(1-\hat{\rho}^{2}\right)-d\sum_{i=1}^{n}\log\left(1+\hat{\rho}^{2}-2\hat{\rho}\bm{y}_{i}^{\top}{\bf m}\right)+\lambda\left({\bf m}^{\top}{\bf m}-1\right).$ (7)
@@
    $\frac{\partial\ell_{SC}}{\partial{\bf m}}=d\sum_{i=1}^{n}\frac{2\rho\bm{y}_{i}}{1+\hat{\rho}^{2}-2\hat{\rho}\bm{y}_{i}^{\top}{\bf m}}+2\lambda{\bf m}={\bf 0}.$
    The updated mean direction is given by the unit vector parallel to $\sum_{i=1}^{n}\frac{\hat{\rho}\bm{y}_{i}}{1+\hat{\rho}^{2}-2\hat{\rho}\bm{y}_{i}^{\top}\hat{{\bf m}}}$.
@@ 2.3 Log-likelihood ratio test for equality of two location parameters
$\ell_{0}=n_{1}d\log\left(1-\rho_{1}^{2}\right)-d\sum_{i=1}^{n_{1}}\log\left(1+\rho_{1}^{2}-2\rho_{1}\bm{y}_{1i}^{\top}{\bf m}_{c}\right)+n_{2}d\log\left(1-\rho_{2}^{2}\right)-d\sum_{i=1}^{n_{2}}\log\left(1+\rho_{2}^{2}-2\rho_{2}\bm{y}_{2i}^{\top}{\bf m}_{c}\right),$ (9)
where $n_{j}$ and $\bm{y}_{ji}$ refer to the $j$-th sample size and the $i$-th observation of sample $j$, respectively, for $j=1,2$. Under the alternative hypothesis, the two location parameters are not equal and hence the relevant log-likelihood to be maximised is given by
$\ell_{1}=n_{1}d\log\left(1-\rho_{1}^{2}\right)-d\sum_{i=1}^{n_{1}}\log\left(1+\rho_{1}^{2}-2\rho_{1}\bm{y}_{1i}^{\top}{\bf m}_{1}\right)+n_{2}d\log\left(1-\rho_{2}^{2}\right)-d\sum_{i=1}^{n_{2}}\log\left(1+\rho_{2}^{2}-2\rho_{2}\bm{y}_{2i}^{\top}{\bf m}_{2}\right),$ (10)
@@ 2.4 Regression analysis
When a set of $p$ covariates is present, we link the location parameter to the covariates by $\bm{\mu}_{i}={\bf B}^{\top}{\bm{x}}_{i}$, where ${\bf B}=\left(\bm{\beta}_{1},\ldots,\bm{\beta}_{d+1}\right)$ denotes the matrix of the regression coefficients, $\bm{\beta}_{i}\in\mathbb{R}^{p}$, and ${\bm{x}}_{i}\in\mathbb{R}^{p}$ denotes a column of the design matrix. The relevant log-likelihood, using (2), becomes
$\ell_{SC}=n\log\left(C_{d}\right)-d\sum_{i=1}^{n}\log\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right),$ (11)
where $\alpha_{i}=\bm{y}_{i}^{\top}\bm{\mu}_{i}$ and $\gamma_{i}=\|\bm{\mu}_{i}\|$.
The advantage of the parametrization in Eq. (2) is evident in the regression setting. Similarly to Paine et al., (2020) and Tsagris and Alzeley, (2024) the errors are heteroscedastic, because we do not assume a common concentration parameter. Each directional vector has its own concentration parameter $\gamma_{i}$ that is linked to $\rho_{i}$ as mentioned earlier. The matrix of the regression coefficients is again estimated via the NR algorithm.
The asymptotic distribution of the maximum likelihood estimator for this regression model is available. Let $\bm{\beta}=(\bm{\beta}_{1}^{\top},\ldots,\bm{\beta}_{d+1}^{\top})^{\top}$ denote the $(d+1)p$-dimensional parameter vector obtained by vectorizing the matrix $\bf{B}$. Then, under certain regularity conditions, the following holds for the maximum likelihood estimator $\hat{\bm{\beta}}_{SC}$ of $\bm{\beta}$:
$\sqrt{n}\,{\cal I}_{SC}^{1/2}\left(\hat{\bm{\beta}}_{SC}-\bm{\beta}\right)\stackrel{{\scriptstyle{\rm d}}}{{\longrightarrow}}N\left(\bm{0},{\bf I}_{(d+1)p}\right)\quad(n\rightarrow\infty).$ (12)
Here ${\cal I}_{SC}$ is the observed Fisher information matrix given by
$\begin{split}{\cal I}_{SC}=&\ \frac{d}{n}\sum_{i=1}^{n}\Biggl{[}\frac{{\bf I}_{d+1}\otimes\bm{x}_{i}\otimes\bm{x}_{i}^{\top}-\sqrt{\gamma_{i}^{2}+1}\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)^{\top}}{\sqrt{\gamma_{i}^{2}+1}\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)}\\ &-\frac{\left\{\left(\bm{\mu}_{i}-\sqrt{\gamma_{i}^{2}+1}\,\bm{y}_{i}\right)\otimes\bm{x}_{i}\right\}\left\{\left(\bm{\mu}_{i}-\sqrt{\gamma_{i}^{2}+1}\,\bm{y}_{i}\right)\otimes\bm{x}_{i}\right\}^{\top}}{(\gamma_{i}^{2}+1)\left(\sqrt{\gamma_{i}^{2}+1}-\alpha_{i}\right)^{2}}\Biggr{]},\end{split}$ (13)
where $\otimes$ denotes the Kronecker product.
@@
where $\alpha_{i}$ and $\gamma_{i}$ are the same as in the case of the SC regression. Asymptotic normality also holds for the maximum likelihood estimator of the PKB regression model under regularity conditions. Let $\hat{\bm{\beta}}_{PKB}$ denote the maximum likelihood estimator of $\bm{\beta}$ for the PKB regression. Then, we have
$\sqrt{n}\,{\cal I}_{PKB}^{1/2}\left(\hat{\bm{\beta}}_{PKB}-\bm{\beta}\right)\stackrel{{\scriptstyle{\rm d}}}{{\longrightarrow}}N\left(\bm{0},{\bf I}_{(d+1)p}\right)\quad(n\rightarrow\infty),$ (15)
where the observed Fisher information matrix ${\cal I}_{PKB}$ is
$\begin{split}{\cal I}_{PKB}=\ &\frac{d+1}{2d}{\cal I}_{SC}+\frac{d-1}{2n}\sum_{i=1}^{n}\Biggl{[}\frac{\sqrt{\gamma_{i}^{2}+1}-1}{\gamma_{i}^{2}}{\bf I}_{d+1}\otimes\bm{x}_{i}^{\top}\otimes\bm{x}_{i}\\ &+\frac{3\gamma_{i}^{2}-2\sqrt{\gamma_{i}^{2}+1}+2}{\gamma_{i}^{4}(\gamma_{i}^{2}+1)}\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)\left(\bm{\mu}_{i}\otimes\bm{x}_{i}\right)^{\top}\Biggr{]}.\end{split}$ (16)
@@ 2.5 Discriminant analysis
Furthermore, we assume that we have at hand a labeled directional dataset, that is, for each observation $\textbf{y}_{i}\in\mathbb{S}^{d}$, the allocation variable $c_{i}\in\{1,\ldots,J\}$ is also recorded, $i=1,\ldots,n$. The goal is to assign a feature observation $\textbf{y}_{0}$ into one among the $J$ classes. Under the maximum likelihood discriminant analysis framework (Mardia et al.,, 1979, 301–305), the rule is to allocate a new observation vector $\bm{y}_{0}\in\mathbb{S}^{d}$ in the group whose likelihood is maximized, i.e,
@@
$\log{\left(\sqrt{\|\boldsymbol{\mu}_{j}\|^{2}+1}-\bm{y}^{\top}_{0}\boldsymbol{\mu}_{j}\right)},\ \ j=1,\ldots,J$
and allocate $\bm{y}_{0}$ to the group $j$ that minimizes the above quantity. In practice, these parameters are unknown, so they are estimated, separately, for each group. In the case of the PKB distribution the allocation rule is straightforward to write down analytically, but evidently, in a more complicated form. In this case, $\bm{y}_{0}$ is allocated to the group $j$ with the minimum value in the following quantity
@@ 2.6 Model based clustering
where $f(\cdot;\bm{\mu}_{j})$ denotes the probability density function of the SC or PKB distribution with (unconstrained) parameter $\bm{\mu}_{j}\in\mathbb{R}^{d+1}$ for $j=1,\ldots,K$. Notice that the previous equation in case where the $j$-th element of $\bm{Z}_{i}$ is equal to 1, collapses to $\bm{Y}_{i}|(Z_{ij}=1)\sim f(\cdot;\bm{\mu}_{j})$. It follows that the joint distribution of $\bm{Y}_{i},\bm{Z}_{i}$ is
@@
Initialization of the EM algorithm demands extra care in order to avoid convergence to local maxima. Some indicative works on this topic include Biernacki et al., (2003), Karlis and Xekalaki, (2003), Fraley et al., (2005), Baudry and Celeux, (2015), Papastamoulis et al., (2016), Michael and Melnykov, (2016). In our implementation of the mixtures of SC distributions, the suggested initialization is based on $k$-means starting values since we observed that for this particular class of finite mixture models this strategy gave comparable results to more sophisticated initialization schemes (see Section 2.3 of Papastamoulis, (2023)). For the case of the PKB we relied on the implementation of Golzy and Markatou, (2020).
@@ 3.2 Hypothesis testing for two location parameters
Table 3: Estimated type I error (nominal error is set to 5%) and estimated power when the data are generated from the SC and PKB distributions with concentration parameters equal to $\rho_{1}=0.3$ and $\rho_{2}=0.8$, for the first and second sample, respectively. The angular difference between the two location directions is given by $\theta$.